Hello! Please help me with this, I was able to answer the other items but this one i was having a hard time on. If you could also show how you solved it i'll really appreciate it, thank you
1) 7x+y=-24
-7x-2y=22
my answer is (-26/7,2) but im not sure, please correct me
2) 2x-6y=4
x+6y=5
my answer is (3,2/6) but im not sure, please correct me
3) 2x+y=1
10x-4y=2
4) y=3x-1
2x-2y=-6

Answer by Edwin McCravy(20054)

Hello! Please help me with this, I was able to answer the other items but this one i was having a hard time on. If you could also show how you solved it i'll really appreciate it, thank you
1) 7x+y=-24
-7x-2y=22
my answer is (-26/7,2) but im not sure, please correct me
That's correct!
2) 2x-6y=4
x+6y=5
my answer is (3,2/6) but im not sure, please correct me
That's correct except for the fact that 2/6 reduces to 1/3
3) 2x+y=1
10x-4y=2
Answer (1/3,1/3)
4) y=3x-1
2x-2y=-6
Answer (2,6)

Edwin